# SECURE: Benchmarking Generative Large Language Models for Cybersecurity Advisory

### v1

@@ 5.10. Ethical concerns
All of the evaluation tasks in our proposed benchmark SECURE use publicly available threat information from credible sources like MITRE (mit, 2024c) (mit, 2024a), CVE (Project, 2024a), CWE (Project, 2024b), and Cybersecurity and Infrastructure Security Agency (CISA) (Advisories, 2024). None of the datasets contain any personal information, and they do not make sensitive judgments on social issues for bias, deception, or discrimination.
6. Limitations & Future Work
@@
We focused on three distinct knowledge tasks (extraction, understanding, and reasoning) to assess the knowledge acquired by LLMs in the cybersecurity context, especially for ICS. While our current work involved designing six datasets centered on ICS security, we plan to extend this framework to encompass other areas of the cybersecurity industry. By adding more datasets, we aim to increase the evaluation “breadth” and provide a more comprehensive assessment. In future work, we also intend to explore various dimensions of trustworthiness, including toxicity, bias, adversarial robustness, out-of-distribution robustness, privacy, and fairness. These investigations are expected to reveal unidentified vulnerabilities and threats to the reliability of using LLMs in cybersecurity. Looking beyond the evaluation of existing LLMs, our research proposes an intriguing direction for enhancing their cybersecurity capabilities. We plan to explore techniques such as Retrieval Augmented Generation (RAG) to fine-tune large language models (Lewis et al., 2020). By training these models on curated cybersecurity corpora, we aim to significantly improve their performance on tasks such as vulnerability identification, threat analysis, and security policy generation. This approach holds immense promise for developing more specialized and effective LLMs tailored to meet the complex demands of the cybersecurity domain.
7. Conclusion
@@
Large Language Models (LLMs) are rapidly transforming the landscape of generative AI, with vast potential applications in cybersecurity. However, concerns regarding their transparency, reliability, and comprehensive understanding, particularly within the sensitive domain of cybersecurity, remain paramount. The SECURE benchmark, introduced in this paper, aims to address these concerns by providing a holistic framework for evaluating the knowledge and capabilities of LLMs in a cybersecurity advisory context. Our rigorous experiments and analysis demonstrate the potential of LLMs as valuable tools for cybersecurity professionals. However, our findings also underscore the need for caution, especially when handling critical or sensitive information. By open-sourcing our benchmark datasets, we invite the research community to contribute to their improvement, enhancing LLM reliability and paving the way for responsible and beneficial generative AI in cybersecurity.
Appendix
Appendix A Large Language Models
Lage Language Models (LLMs) are pretrained autoregressive transformers (Vaswani et al., 2017) capable of generating one token at a time given a set of generated tokens as input. These are trained to maximize the probability of the next token given a history of tokens. We use the following LLMs in our evaluation:
  1. (1)
    ChatGPT models: ChatGPT models by OpenAI have brought remarkable improvements in LLMs and sparked a trend for building more intelligent AI models. We use the following LLMs from OpenAI in our work, accessed via OpenAI’s API querying system (gpt, 2024d):
    a) GPT-3.5 models can understand and generate natural language. We use the gpt-3.5-turbo model that has been optimized for chat. It has a context window of 16385 tokens and the model was trained up to the cutoff time of September 2021.
    b) GPT-4 is a multimodal modal that can accept both text and image as inputs. This model excels in solving various linguistic tasks compared to GPT-3.5 model. We use the gpt-4-turbo version which is optimized for chat. It has a context window of 128000 tokens. The model was trained on a dataset collected upto to December 2023.
  2. (2)
    Llama3-70B and Llama3-8B: Llama3 models (lla, 2024a) are a family of large language models developed by Meta, which is open-sourced to the researchers. These models are available in parameter sizes of 8B and 70B with both pre-trained and instruction-tuned (optimized for chat) available in public. These models input text only, and generate text and code only. We use these models through huggingface API. Llama3 models were released on April 18, 2024.
  3. (3)
    Gemini (gemmini): Gemini models are a family of multimodal generative models released by Google which are capable of handling extremely long contexts, and take input as documents, video, audio and pictures. We use the Gemini-Pro 1.5 version of the LLM available through API.
  4. (4)
    Mistral-7B and Mixtral 8x7B: Mixtral 8x7b (mix, 2024) is a mixture-of-experts model ( Mixtral-7B (Jiang et al., 2023)), consisting of 8 distinct groups of parameters. We use both of these models through huggingface API.
Appendix B Prompt templates for evaluation
B.1. Extraction Task
Example Prompt (Task: MAET)
For the given question: Which protocol function can be disabled to prevent unauthorized device shutdowns?, and four options: A) DNP3 0x0D, B) HTTP GET, C) SMTP HELO, or D) FTP LIST, pick the best option as the answer, and return as either A, B, C or D. If you do not know the answer, return X. Choose the appropriate letter from A, B, C, D, or X as your answer. Please provide only the letter corresponding to your choice without any additional text or explanations.
Example Prompt (Task: CWET)
For the given question: What is a common method used by attackers to bypass ATA password security?, and four options: A) Using a BIOS exploit, B) Hot swapping the drive, C) Encrypting the drive, or D) Using default passwords, pick the best option as the answer, and return as either A, B, C or D. If you do not know the answer, return X. Choose the appropriate letter from A, B, C, D, or X as your answer. Please provide only the letter corresponding to your choice without any additional text or explanations.
B.2. Understanding Task
Example Prompt (Task: KCV)
You are given the following JSON data as context: $JSON\_FILE$. Based on the context, you have to analyze the following statement: PAN-OS versions 9.0.17 and later are unaffected by the vulnerability. and indicate whether the statement is True or False. Return your answer as either T (for True) or F (for False). If you do not know the answer, return X. Provide only the letter corresponding to your choice (T, F, or X) without any additional text or explanations.
For this example we pass the $JSON\_file$ downloaded from this CVE-Link [^2] https://github.com/CVEProject/cvelistV5/blob/main/cves/2024/0xxx/CVE-2024-0008.json.
Example Prompt (Task: VOOD) In VOOD, we do not pass any context to the LLMs and inspect how the models perform on information that has not been seen by them during training.
For the given statement: The vulnerability described in CVE-2024-0011 allows for the execution of arbitrary code on the affected system., indicate whether the statement is True or False. Return your answer as either T (for True) or F (for False). If you do not know the answer, return X. Provide only the letter corresponding to your choice (T, F, or X) without any additional text or explanations.
B.3. Reasoning Task
Example Prompt (Task: RERT)
You are given the following descriptions of a cyber incident Vulnerability Overview. Please perform an analysis and deduce the risk evaluation. Return your answer in one sentence in a format such as ”Successful exploitation of this vulnerability could allow an attacker to inject arbitrary JavaScript into a user’s web browser” for a single vulnerability, or ”Successful exploitation of these vulnerabilities could cause a denial of service, disclosure of sensitive information, communication loss, and modification of settings or ladder logic” for multiple vulnerabilities. 3.2.1 Path Traversal CWE-22 There are multiple ways in LAquis SCADA for an attacker to access locations outside of their own directory. CVE-2024-5040 has been assigned to this vulnerability. A CVSS v3.1 base score of 7.8 has been calculated; the CVSS vector string is (AV:L/AC:L/PR:N/UI:R/S:U/C:H/I:H/A:H). A CVSS v4 score has also been calculated for CVE-2024-5040. A base score of 8.5 has been calculated; the CVSS vector string is (CVSS4.0/AV:L/AC:L/AT:N/PR:N/ UI:P/VC:H/VI:H/VA:H/SC:N/SI:N/SA:N).
Example Prompt (Task: CPST)
Calculate the CVSS v3 score for the following CVSS v3 Vector String: AV:L/AC:L/PR:N/UI:R/S:U/C:H/I:H/A:H. Return your answer as a single number between 0.0 and 10.0. Provide only the number corresponding to your answer without any additional text or explanations.
Appendix C LLM Reasoning Example
Example LLAMA3-70b reasoning for one question in the KCV task
To determine the accuracy of the statement, I analyzed the provided JSON data, which describes a CVE record for CVE-2024-0011. The vulnerability is a reflected cross-site scripting (XSS) issue in the Captive Portal feature of Palo Alto Networks PAN-OS software. The description of the vulnerability states: ”A reflected cross-site scripting (XSS) vulnerability in the Captive Portal feature of Palo Alto Networks PAN-OS software enables execution of malicious JavaScript (in the context of an authenticated Captive Portal user’s browser) if a user clicks on a malicious link, allowing phishing attacks that could lead to credential theft.” The key points to note are: 1. The vulnerability is a reflected XSS issue, which means it allows an attacker to inject malicious JavaScript code into a user’s browser. 2. The vulnerability is limited to the context of an authenticated Captive Portal user’s browser. 3. The goal of the attack is to perform phishing attacks that could lead to credential theft. Nowhere in the description or the provided data is there any indication that the vulnerability allows for the execution of arbitrary code on the affected system. The vulnerability is primarily a client-side issue, affecting the user’s browser, rather than a server-side issue that could lead to arbitrary code execution on the system. Therefore, based on the analysis of the provided data, I conclude that the statement is False.

### v3

@@ V-K Ethical concerns
All of the evaluation tasks in our proposed benchmark SECURE use publicly available threat information from credible sources like MITRE [11] [12], CVE [13], CWE [14], and Cybersecurity and Infrastructure Security Agency (CISA) [15]. None of the datasets contain any personal information, and they do not make sensitive judgments on social issues for bias, deception, or discrimination.
VI Customized LLMs
VI-A Retrieval Augmented Generation
Retrieval Augmented Generation (RAG) is a method that combines the strengths of retrieval based models and generation based models [58]. It consists of two major components: retriever and generator. The retriever searches for relevant information from a large database or knowledge base, using retrieval techniques to find the most relevant documents, paragraphs, or data points. Once relevant information is retrieved, the generator component (typically a language model like GPT) uses this information as context to generate a more accurate and contextually relevant response.
Figure 11: RAG Model
We use LLAMA3-8B as our generator model and design retriever to search for relevant information from sources like MITRE, CWE, and CVEs. We utilize the LangChain framework [59] for data-handling. The documents are segmented into chunks of 512 tokens with an overlap of 20 tokens using LangChain’s RecursiveCharacterTextSplitter. This chunk size ensures that the model retains enough contextual information. The chunks are then transformed into vector embeddings using the ”mixedbread-ai/mxbai-embed-large-v1” model [60], and stored these embeddings in the FAISS database [61].
When a user queries the RAG model, it is first converted into vector embeddings using the same ”mixedbread-ai/mxbai-embed-large-v1” model. These embeddings are then processed through LangChain’s ConversationalRetrievalChain, which identifies the closest matches to the query from the stored documents. The matched documents, along with the query, are then fed into Llama3 to generate a coherent response. Figure 11 shows the block diagram of our RAG model.
VI-B Fine-tuning
Fine-tuning is the process of taking a pretrained model and training it further on a specific dataset to adapt it to particular tasks. The goal of this ‘tuning’ is to help the model learn the nuances of the new data. We follow the following steps:
Data preparation: We collect data using the web-graph data from Common Crawl [62], and filter nodes of representative domains, based on their eigenvector centrality. We select the top 1 million nodes, ranked by highest centrality, for community detection, and scrap web pages from the identified domains within the community using Trafilatura [63]. We use this dataset for continual pre-training but exclude from the instruction fine-tuning stage.
We also utilize a a set of human-collected URLs to ensure the dataset is specific to Industrial Control Systems (ICS) by sorting and filtering the data extracted with Trafilatura. This step ensures the dataset focuses on ICS-specific content before proceeding to create the instruction dataset.
Instruction generation and fine-tuning: We apply the Self-Instruct method to generate instructions from the parsed text, using the default system prompt from the Self-Instruct paper [64]. This step yielded 10,000 instructions. We generate responses using the Llama3.1 70B Instruct model, with 8-bit quantization, and a custom system prompt. We then apply instruction fine-tuning on the Llama3-8B Instruct model with 4-bit training.
TABLE VII: Benchmark results on customized LLMs
Model MAET CWET KCV
Base-Llama3-8B 82.1% 83.9% 82.8%
RAG model 86.6% 77.3% 60.5%
Fine-tuned model 84% 85% 76%
VI-C Results:
Table VII shows the result of the benchmark. MAET, CWET, and KCV do not have source URLs provided as a context, and hence can accurately evaluate customized models performance against general LLMs. We can observe that both of the customized models perform better on MAET. Fine-tuned model outperform RAG on CWET. While fine-tuned model performs comparable to base model on KCV, RAG model has a significant drop. One key weakness of “Retrieve-Read” framework for RAG is that during the retrieval phase, the model can select irrelevant chunks and miss crucial information for accurate response. Hence, if the content is not supported by the retrieved context, the model faces hallucinations [65]. This could be one of the reasons behind lower performance in CWET and KCV. We will explore advanced RAGs framework in our future works.
VII Limitations & Future Work
The reliance on standardized sources such as MITRE ATT&CK, CWE, CVE, and CVSS introduces potential biases and coverage gaps in our benchmark datasets. While these sources are integral to the cybersecurity community and are frequently used in research, they are not entirely without limitations. In particular, issues such as incorrect versioning in CVE entries or incomplete coverage of attack techniques in MITRE ATT&CK could impact the accuracy of SECURE’s evaluation tasks.
To address this, we have validated the datasets through expert review and cross-referencing with additional sources. However, these measures cannot eliminate the inherent biases of the original databases. In future work, we will continue refining these datasets and exploring alternative data sources.
@@ VII Limitations & Future Work
We focused on three distinct knowledge tasks (extraction, understanding, and reasoning) to assess the knowledge acquired by LLMs in the cybersecurity context, especially for ICS. While our current work involved designing six datasets centered on ICS security, we plan to extend this framework to encompass other areas of the cybersecurity industry. By adding more datasets, we aim to increase the evaluation “breadth” and provide a more comprehensive assessment. In future work, we also intend to explore various dimensions of trustworthiness, including toxicity, bias, adversarial robustness, out-of-distribution robustness, privacy, and fairness. These investigations are expected to reveal unidentified vulnerabilities and threats to the reliability of using LLMs in cybersecurity.
VIII Conclusion
@@
LLMs are rapidly transforming the landscape of AI, with vast potential applications in cybersecurity. However, concerns regarding their reliability, and understanding, particularly within the sensitive domain of cybersecurity, remain paramount. The SECURE benchmark, introduced in this paper, aims to address these concerns by providing a holistic framework for evaluating the capabilities of LLMs in a cybersecurity advisory context. Our experiments and analysis demonstrate the potential of LLMs as valuable tools for cybersecurity professionals. However, our findings also underscore the need for caution, especially when handling reasoning tasks. By open-sourcing our benchmark datasets, we invite the research community to contribute to their improvement, enhancing LLMs reliability and paving the way for responsible and beneficial large language models in cybersecurity.
Acknowledgement
This work is supported by Grant H98230-21-1-0317.